Consider how the methods of undetermined coefficients and variation of parameters can be combined to solve the given differential equation. Carry out your ideas to solve the differential equation.

y'' − 2y' + y = 4x^2 − 1 + x^−1e^x

y(x) =____
1
Expert's answer
2020-07-07T20:35:20-0400

Given differential equation is "y'' -2y' +y = 4x^2 - 1 + \frac{e^x}{x}"

Auxiliary equation is "m^2 - 2m + 1 = 0 \implies m =1,1"

so "y = C_1e^x+C_2xe^x"


Let us solve for ("4x^2-1)" by undetermined coefficients and for "\frac{e^x}{x}" by variation of parameters


(a) undetermined coefficients:

let "y = Ax^2 +Bx+C" be solution of the equation "y'' -2y' +y = 4x^2 - 1"

then differentiating y first and second time with respect to x, and putting in differential equation, then equation will be ,

"2A-4Ax-2B+Ax^2 +Bx+C = 4x^2-1 \implies A = 4, B=16, C=23"

then "y = C_1e^x+C_2xe^x+ 4x^2 +16x+23"


(b) method of variation of parameters

Differential equation is "y'' -2y' +y = \frac{e^x}{x}"

since "y = C_1e^x+C_2xe^x" for auxiliary equation

Here "u = e^x, v=xe^x, R =\frac{e^x}{x}"

Then W, Wronskian of u and v will be


"W = \begin{vmatrix}
 e^x & xe^x \\
 e^x & e^x(x+1)
\end{vmatrix} = e^{2x}"


then "P.I. = -u\int\frac{vR}{W}dx + v\int\frac{uR}{W}dx"


"P.I. = e^x\int\frac{xe^xe^x}{xe^{2x}}dx + xe^x\int\frac{e^{2x}}{xe^{2x}}dx"


"P.I. = -e^x\int dx + xe^x\int \frac{dx}{x} = xe^x(ln(x) - 1)"



Hence, final solution of the equation will be


"y = C_1e^x+C_2xe^x+ 4x^2 +16x+23 + xe^x(ln(x)-1)"